1. Lelentle has purchased 10 of the 1000 tickets sold in a raffle. To determine the 5 prize winners, 5 tickets are drawn at random. Compute the probability that Lelentle will win at least 1 prize.
1
Expert's answer
2020-10-02T14:51:38-0400

Let p be the probability that they do not win the prize and 1-p be the probability of winning the prize. As it is given in 1000 ticket ,10 ticket sold as we have to determine probability if 5 ticket drawn at random and in that 1 ticket selected randomly

so here ticket that dont give prize

"p=\frac{^{995}C_{10}}{^{1000C_{10}}}"


so probability of getting prize = 1-p

P=1- "\frac{^{995}C_{10}}{^{1000C_{10}}}"

P=0.05 is the required probability